OpenStudy (anonymous):

The product of two consecutive whole numbers is 1.5 times greater than the square of the smaller number. Find the two numbers

OpenStudy (anonymous):

@ghazi

OpenStudy (anonymous):

x+(x+1) = 1,5/(x^2) Find x with the equation and you'll know y as well..

OpenStudy (whpalmer4):

That should be \[x*(x+1) = 1.5x^2\]

OpenStudy (anonymous):

@whpalmer4 I have two answers for x = 0 and 2

OpenStudy (whpalmer4):

That's what I get as well.